How many gallons of 25% antifrezze solution should be added to 5 gallons of a 10% solution to obtain a 15% solution?
Answer by mananth(16946)

percent ---------------- quantity
Antifreeze I 10 ---------------- 5 gal
Antifreeze II 25 ---------------- x gal
Mixture 15 ---------------- 5 + x gal

10*5+25x=15(5+x)
50 + 25 x = 75 + 15 x
25 x + -15 x = 75 - -50
10 x = 25
/ 10
x = 2.5
2.5 gal of 25 % antifreeze II